radioactive substance has half life of 1 hour . therefore if two nuclei of the substance are present initially after one hour only one will remain disassociated.
 
This statement is false. but what is the reason for this statement to be false ?

Grade:12th pass

1 Answers

Mahima Kanawat
5 years ago
Dear student.... Radioactive reactions follow first order kinetics, so, N =N°(1/2)^t/T ; N=2(1/2)^1/1 ; N= 1, so this statement is true that after 1 hour 1 nuclei will remain undissociated.... WITH REGARDS, MAHIMA (ASKIITIANS FORUM EXPERT)